OpenStudy (anonymous):

determine the nature of the solution to the equation 4x^2=3x-5

OpenStudy (amistre64):

nature?

OpenStudy (amistre64):

sounds like they want to know what the discriminant will produce

OpenStudy (anonymous):

A two irrational solutions B. two complete solutions that are not real C. one repeated real solution D. two radical solutions

OpenStudy (anonymous):

check disciminant value : 4x^2=3x-5 4x^2-3x+5 = 0 D=b^2-4ac D=(-3)^2-4*4*5 D=9-80 D=-71 because D<0, kind of roots is not real